### v1

@@ Abstract
In recent years, federated learning (FL) has emerged as a promising approach for machine learning (ML) and data science across distributed edge devices. With the increasing popularity of FL, resource contention between multiple FL jobs training on the same device population is increasing as well. Scheduling edge resources among multiple FL jobs is different from GPU scheduling for cloud ML because of the ephemeral nature and planetary scale of participating devices as well as the overlapping resource requirements of diverse FL jobs. Existing resource managers for FL jobs opt for random assignment of devices to FL jobs for simplicity and scalability, which leads to poor performance.
In this paper, we present Venn, an FL resource manager, that efficiently schedules ephemeral, heterogeneous devices among many FL jobs, with the goal of reducing their average job completion time (JCT). Venn formulates the Intersection Resource Scheduling (IRS) problem to identify complex resource contention among multiple FL jobs. Then, Venn proposes a contention-aware scheduling heuristic to minimize the average scheduling delay. Furthermore, it proposes a resource-aware device-to-job matching heuristic that focuses on optimizing response collection time by mitigating stragglers. Our evaluation shows that, compared to the state-of-the-art FL resource managers, Venn improves the average JCT by up to 1.88×1.$8$.
footnotetext: 1 Computer Science and Engineering, University of Michigan, Michigan, USA 2 Computer Science and Engineering, University of Illinois at Urbana-Champaign, Illinois, USA. Correspondence to: Jiachen Liu <amberljc@umich.edu>.
@@ 1 Introduction
Federated learning (FL) enables distributed edge devices to perform collaborative machine learning (ML) without moving raw data into the cloud Bonawitz et al. (2019); Ramage & Mazzocchi (2020). FL has been adopted by many large corporations including Apple, Meta, Google, and LinkedIn to protect user data privacy while improving user experience. For example, Google adopts FL for a wide range of applications such as speech recognition Warden (2018), healthcare study Sadilek et al. (2021), next-word prediction Hard et al. (2019); Xu et al. (2023), emoji prediction Ramaswamy et al. (2019), and query suggestion on keyboard Yang et al. (2018). Each FL training job in practice often requires 1000$\sim$10000 device participants in each training round and takes 4$\sim$8 days to finish Yang et al. (2018). As the number of FL applications continues to grow, efficient edge resource management has become the key to fast and resource-efficient FL.
As opposed to traditional ML training in the cloud Gu et al. (2019); Xiao et al. (2018); Narayanan et al. (2020), resource management in FL introduces several unique characteristics. First and foremost, the availability of FL resources varies rapidly and unpredictably during the lifetime of an FL job, because client devices may become available, join training, or drop off at any time. Meaning, in addition to job-level unpredictability (e.g., when a training job will converge and finish), we must mitigate device-level unpredictability too.
@@
Second, in comparison to cloud ML, FL involves orders-of-magnitude larger resource pools that are also highly heterogeneous in terms of their hardware capacity, software versions, and training data availability. Even when running on the same population, different FL jobs often compete for different subsets of devices, influenced by their specific model characteristics and training objectives. As such, complex multi-resource contentions are common among multiple ongoing FL jobs, where eligible resources for one job may overlap, contain, or be within those of one or more other jobs.
Figure 1: Composition of the completion time of one round of an FL job.
Finally, due to the collaborative and distributed nature of FL, the job completion time (JCT) of an FL job is significantly affected by both scheduling delay and response collection time (Figure 1). During each training round, the former refers to the time needed to acquire all resources required by the FL job. In contrast, the latter refers to the time needed to collect sufficient responses so that the round can be considered complete. A joint optimization of both these factors is necessary to optimize an FL job’s JCT, but it is inherently challenging due to the trade-off between the two objectives. On the one hand, reducing response collection time often involves cherry-picking ideal devices based on additional criteria, which could prolong the scheduling delay of the job. On the other hand, coarse-grained client selection that promptly assigns all incoming eligible devices prevents swift and qualified responses, which eventually takes more time to collect enough satisfactory responses. Importantly, jointly optimizing job scheduling and job-to-device matching is a well-known NP-hard problem Gu et al. (2019), and it is further magnified by the scale of the devices and the FL-specific contention.
@@
In the early days of FL research Kairouz et al. (2019), most solutions focused on optimizing the response collection time of individual jobs, while ignoring the scheduling delay. With the proliferation of FL in production, large corporations such as Apple Paulik et al. (2021), Meta Huba et al. (2022), and Google Bonawitz et al. (2019) are running many FL jobs concurrently. As a result, they have developed their own FL infrastructures to coordinate multiple FL jobs at production scale. However, despite low-level differences, these FL resource managers can all be described simply as random device-to-job matching in various forms. We show that such random assignment can lead to higher scheduling delays and response collection times as more and more FL jobs vie for a shared device population. Note that prior works on resource allocation with job-specific constraints Thinakaran et al. (2017) are not suitable here either, because they were designed for datacenter resources that are not ephemeral, nor so numerous.
@@
In this paper, we present Venn, an FL resource manager that minimizes the average JCT of multiple FL jobs competing over a large number of heterogeneous devices in two steps. First, Venn addresses the intricate resource contention among FL jobs by formulating it as an Intersection Resource Scheduling (IRS) problem (§4.2), where the resources that one job contends for may overlap, contain, or be within those resources of other jobs. Venn then introduces a contention-aware scheduling heuristic that prioritizes small jobs requiring scarce resources to minimize the average scheduling delay. Second, Venn employs a resource-aware device-to-job matching heuristic to reduce response collection time (§4.3). Together, they jointly optimize the average JCT for a diverse set of FL jobs operating under dynamically changing and uncertain resource conditions.
@@
We have implemented and evaluated Venn across various FL workloads derived from real-world scenarios (§5). Compared to state-of-the-art FL resource allocation solutions Bonawitz et al. (2019); Huba et al. (2022); Paulik et al. (2021), Venn improves the average JCT by up to 1.88×1.$8$.
@@
    We introduce Venn, an FL resource manager designed to enable efficient sharing of heterogeneous devices across a large number of FL jobs.
@@
    To minimize the average JCT of FL jobs, we propose a scheduling and matching joint solution to optimize both the scheduling delay and response collection time.
@@
    We have implemented and evaluated Venn, along with its scheduling and matching algorithms, demonstrating improvements in the average JCT compared to the state-of-the-art across various real-world FL workloads.
@@ 2 Background and Motivation
2.1 Federated Learning
FL has been widely adopted by the industry to train ML models on larger and more diverse on-device datasets without having to copy raw data into the cloud. Example applications include healthcare study, speech recognition, next-word prediction, etc.. FL is unique both in terms of resources it uses and jobs that use it, which pose unique challenges for FL resource managers.
@@ 2.1 Federated Learning
Figure 2: FL resources exhibit both high variance in availability and capacity.
@@
(b) Random Matching (6.33).
(c) SRSF (5).
(d) Optimal (4.33).
Figure 3: Toy example of resource allocation for three schedules (with tie-breaking based on job ID) across jobs with varying resource demands and device eligibility requirements. Numbers in parentheses are the average JCTs. Devices check in at a constant rate. Eligible devices only for Emoji jobs are marked with blue; all devices are eligible for the Keyboard job. The label of each client indicates its job assignment. Random Matching and SRSF inefficiently allocate scarce Emoji-eligible devices to Job 1, which already has ample Keyboard-eligible resources. Conversely, the optimal schedule smartly allocates these scarce resources to Jobs 2 followed by Job 3, thereby minimizing average request completion time.
FL Resources
FL resources often refer to edge devices with limited capacities, including smartphones, laptops, and Internet of Things (IoT) devices. Compared to resources used by traditional ML cluster (e.g., GPUs), FL resources are unique in terms of availability and heterogeneity.
Dynamic availability. As FL moves ML computations to user devices, the availability of edge devices for sufficiently long durations become crucial. We analyze a real-world client availability trace Lai et al. (2021a), which encompass 180 million trace items of devices behavior over a week. Figure 1(a) shows that the number of available devices (charging and connected to WiFi) changes over time with diurnal pattern. Hence, there is an inherent scheduling delay for an FL job to acquire a sufficient number of device resources.
Device heterogeneity. The large number of devices involved in FL are heterogeneous in multiple dimensions including data availability, software version, and hardware capacity, which leads to varying response time among devices. Figure 1(b) showcases this heterogeneity, focusing on variations in memory and CPU capacities among edge devices, based on data from Ignatov et al. (2019). It also annotates the minimum hardware requirements needed to execute three popular on-device ML models within a reasonable time. As jobs may have different learning objectives, they may compete for different subsets of devices. When more factors like data availability and software version are considered, these subsets can exhibit relationships that are inclusive, overlapping, or nested, leading to complex contention patterns.
FL Jobs
An FL job is executed in multiple rounds that run sequentially for synchronous FL training. Each round begins by issuing a resource request including its demand size and device specifications. Once the request is sent, that particular round of the job enters a waiting phase where it continuously dispatches tasks to participants and collects their responses until a sufficient number of qualified responses are gathered. This process repeats for each round until all the rounds and therefore the job completes (e.g., it reaches a target accuracy or completes a pre-defined number of rounds).
In this paper, we focus only on synchronous FL jobs.
2.2 FL Resource Managers
To orchestrate large FL training jobs, several FL resource managers have been proposed in the industry with three primary designs that all support synchronous FL jobs.
@@ 2.2 FL Resource Managers
    Apple’s FL resource manager Paulik et al. (2021) is driven by clients, where each client independently samples from a list of FL jobs they are able to execute.
@@
    Meta’s FL resource manager Huba et al. (2022) is centralized, where it randomly matches each client with one eligible FL job.
@@
    Google’s FL resource manager Bonawitz et al. (2019) is driven by jobs, where each job independently samples from available clients.
Despite the seeming variety in their designs, existing resource managers all boil down to random device-to-job matching in different forms. Unfortunately, in the absence of extreme surplus of qualified devices compared to the demand, random matching fail to address the growing resource contention and yield longer average scheduling delay.
@@
Consider the toy example in Figure 3 that compares three centralized scheduling solutions: random matching as specified above, Shortest-Remaining-Service-First (SRSF) proposed for cloud ML scheduling Gu et al. (2019), and the optimal solution, assuming all jobs arrive simultaneously. Both random matching and SRSF fail to account for resource scarcity and contention, thereby allocating limited resources (i.e., Emoji clients) to jobs (i.e., the Keyboard job) that already have an abundance of resources, resulting in an average JCT of 6.33 and 5 time units. Note that the contention patterns in real-world scenarios are often more complex and larger in scale than the one presented in this example.
2.3 Opportunities for Improvements
Figure 4: Impact of resource contention.
@@ 2.3 Opportunities for Improvements
Impact of resource contention
Most of existing FL solutions that focus on single-job optimizations, such as optimizing for client selection Lai et al. (2021b); Abdelmoniem et al. (2023); Balakrishnan et al. (2022), often rely on the assumption that FL jobs have access to sufficient number of online devices. However, in practice, there could be multiple FL jobs running at the same time and competing for the same set of devices, leading to resource contention Bonawitz et al. (2019). This can significantly impact the performance of FL jobs, such as accuracy and wall clock time.
We analyze the impact of resource contention on FL jobs’ performance in Figure 5. In this experiment, the resource pool is evenly partitioned and managed by each job, who aims to train a ResNet-18 model for the FEMNIST dataset Cohen et al. (2017). We vary the number of concurrently running jobs to observe its impact. As more jobs share the same device pool, the available device choices for each job become increasingly constrained, leading to a noticeable degradation in the round-to-accuracy performance. Hence, evenly partitioned resource pool for each FL job is no better than a shared device pool in terms of participant diversity.
Breakdown of request completion time
While most existing FL efforts, such as quantization Reisizadeh et al. (2020) and client selection He et al. (2023), predominantly focus on optimizing response collection time—–i.e., the time needed to collect a sufficient number of responses—–they often overlook a critical component: scheduling delay, as depicted in Figure 1.
To offer a holistic view of JCT, we dissect its components for a single round request, under varying degrees of resource contention, as shown in Figure 5. Utilizing the same experimental setup as in Figure 5, we quantify both the average scheduling delay and response collection time during one round of training and testing, involving 100 participants. The scheduling delay is specifically captured under random device-to-job matching, and the shaded regions cover the duration for each job. As our results in Figure 5 indicate, scheduling delay can significantly impact overall JCT, especially when resource supply falls short of demand.
@@ 3 Venn Overview
Venn serves as a standalone FL resource manager that operates at a layer above all FL jobs, and it is responsible for allocating each checked-in resource to individual jobs. Figure 6 provides an overview of Venn’s workflow, along with its role in an FL job’s lifecycle.
@@
In each execution round, an FL job submits resource requests to Venn, specifying its device requirements and associated resource demands (0⃝). Devices continuously check in with Venn as they become available over time (1⃝). Based on the real-time resource demand and supply, Venn generates a resource allocation plan to assign one FL job to each checked-in device (2⃝), until the job’s demand is satisfied. Upon receiving the task assignment from Venn, each device adheres to the allocation plan and participates in the corresponding job (3⃝). The device then retrieves the computation plan from the job and performs on-device computation (Lv et al., 2022; CoreML, ) (3⃝ and 4⃝). Finally, the device may either report the training result to the corresponding job upon completion, or drop off mid-process due to availability dynamics (5⃝).
Note that Steps 3⃝ through 5⃝ adhere to conventional FL protocols between individual FL jobs and devices. Venn’s primary role is in optimizing the job-to-device assigning phase, denoted by 2⃝.
Venn delegates responsibilities such as device selection, device fault tolerance, and privacy protection to individual FL jobs. Device failures are both inevitable and difficult to predict in FL. Rather than imposing a one-size-fits-all solution, Venn empowers FL jobs to take the reins on fault tolerance based on their specific workloads and objectives. Therefore, Venn offloads handling device fault tolerance to FL jobs, who can better detect and react to device failures (e.g., deciding the amount of overcommit Bonawitz et al. (2019)). Similarly, Venn offers FL jobs the freedom to design their own device selectors Lai et al. (2021b), where they can incorporate customized resource criteria into their requests. Venn also does not interfere with job-specific privacy solutions such as secure aggregation Bonawitz et al. (2016); Huba et al. (2022) or differential privacy Geyer et al. (2017); Xu et al. (2023).

### v2

@@ Abstract
In recent years, collaborative learning (CL) has emerged as a promising approach for machine learning (ML) and data science across distributed edge devices. As the deployment of CL jobs increases, they inevitably contend for limited resources. However, efficient resource scheduling in this context is challenging because of the ephemeral nature and resource heterogeneity of devices, coupled with the overlapping resource requirements of diverse CL jobs. Existing resource managers often assign devices to CL jobs randomly for simplicity and scalability, but this approach compromises job efficiency.
In this paper, we present Venn, a CL resource manager that efficiently schedules ephemeral, heterogeneous devices among multiple CL jobs to reduce the average job completion time (JCT). Venn formulates the Intersection Resource Scheduling (IRS) problem to identify complex resource contention among multiple CL jobs. It then proposes a contention-aware scheduling heuristic to minimize the average scheduling delay. Furthermore, it proposes a resource-aware device-to-job matching heuristic to optimize response collection time by mitigating stragglers. Our evaluation shows that, compared to the state-of-the-art CL resource managers, Venn improves the average JCT by up to 1.88×1.$8$. The code is available at https://github.com/SymbioticLab/Venn.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
@@ 1 Introduction
Collaborative learning (CL) enables distributed edge devices to perform collaborative machine learning (ML) without moving raw data into the cloud Bonawitz et al. (2019); Ramage & Mazzocchi (2020). CL has been adopted by many large corporations including Apple, Meta, Google, and LinkedIn to protect user data privacy while improving user experience. For example, Google adopts CL for a wide range of applications such as speech recognition Warden (2018), healthcare study Sadilek et al. (2021), next-word prediction Hard et al. (2019); Xu et al. (2023), emoji prediction Ramaswamy et al. (2019), and query suggestion on keyboard Yang et al. (2018). Each CL training job in practice often requires 1000$\sim$10000 device participants in each training round and takes 4$\sim$8 days to finish Yang et al. (2018). As the number of CL applications continues to grow, efficient edge resource management has become the key to fast and resource-efficient CL.
@@
Unlike cloud resources, CL resources are not accessible all the time; they only become available for training when they are connected to WiFi and charging Bonawitz et al. (2019). Moreover, they are highly heterogeneous in terms of hardware capacity, software versions, and training data availability (§2.1). Even when running on the same client population, different CL jobs often compete for different subsets of devices, based on their specific model characteristics and training objectives. The resource heterogeneity and job’s resource requirement diversity lead to complex multi-resource contentions among ongoing CL jobs, where eligible resources for each job are not only limited but also may overlap, contain, or be within those of one or more other jobs.
Figure 1: Composition of the completion time of one round of a CL job.
However, existing CL efforts often overlook resource contention among coexisting CL jobs, assuming that sufficient resources are always available Bonawitz et al. (2019). For example, many recent works Lai et al. (2021a); Abdelmoniem et al. (2023); Lai et al. (2021b); Balakrishnan et al. (2022) primarily focus on optimizing the response collection time of individual CL jobs. In practice, however, multiple CL jobs may run concurrently and compete for a subset of devices, leading to resource contention Bonawitz et al. (2019). As a result, existing works fail to account for scheduling delay (§2.2)—the time needed to acquire all necessary resources for a CL job—as a key factor influencing job completion time (JCT) (Figure 1).
@@
With the proliferation of CL in production, large corporations such as Apple Paulik et al. (2021), Meta Huba et al. (2022), and Google Bonawitz et al. (2019) have developed their own CL infrastructures to manage multiple jobs at scale. However, despite their low-level differences, these CL resource managers can largely be characterized as random device-to-job matching in various forms. We observe that such random matching can result in higher scheduling delays and longer response collection times as an increasing number of CL jobs compete for a shared pool of devices.
@@
In this paper, we present Venn, a CL resource manager designed to minimize the average JCT of multiple CL jobs competing for a large pool of edge resources. First, Venn models the complex resource contention among CL jobs as an instance of the Intersection Resource Scheduling (IRS) problem (§4.2), where a job’s resource demands may overlap with, contain, or fall within those of other jobs. Venn then introduces a contention-aware scheduling heuristic that prioritizes jobs requiring scarce resources and lower overall demand to minimize the average scheduling delay. Second, Venn employs a resource-aware device-to-job matching heuristic to mitigate the effect of stragglers on response collection time (§4.3). Together, these techniques jointly optimize the average JCT of multiple CL jobs operating under dynamically changing and uncertain resource conditions.
@@
We have implemented and evaluated Venn across various CL workloads derived from real-world scenarios (§5). Compared to state-of-the-art CL resource allocation solutions Bonawitz et al. (2019); Huba et al. (2022); Paulik et al. (2021), Venn improves the average JCT by up to 1.88×1.$8$.
@@
    We introduce Venn, a CL resource manager designed to enable efficient sharing of heterogeneous devices across a large number of CL jobs.
@@
    To minimize the average JCT of CL jobs, we propose a job scheduling and client matching joint solution to optimize both the scheduling delay and response collection time. Our approach achieves a logarithmic-linear time complexity and is supported by theoretical analysis.
@@
    We have implemented and evaluated Venn, along with its scheduling and matching algorithms, demonstrating improvements in the average JCT compared to the state-of-the-art across various real-world CL workloads.
@@ 2 Background and Motivation
2.1 Collaborative Learning
Collaborative learning (CL) has been widely adopted in industry to train ML models on diverse datasets stored on edge devices without transferring raw data to the cloud. Example applications include healthcare study, speech recognition, next-word prediction, etc. However, CL introduces unique resource management challenges due to the distinct characteristics of edge devices and the jobs that rely on them.
@@ 2.1 Collaborative Learning
Figure 2: CL resources exhibit both high variance in availability and capacity.
Single CL Job’s Resource Management:
CL resources often refer to edge devices with limited capacities, including smartphones, laptops, and Internet of Things (IoT) devices. A major challenge of resource management in CL stems from the unique characteristics of CL resources in terms of availability and heterogeneity.
Dynamic availability. Unlike cloud resources (e.g., GPUs) that are accessible on demand, edge devices in CL are only available for training under specific conditions, such as when they are charging and connected to WiFi. We analyze a real-world client availability trace Lai et al. (2021a), which encompass 180 million trace items of devices behavior over a week. Figure 2(a) shows that the number of available devices (charging and connected to WiFi) changes over time with diurnal pattern. This fluctuation creates inherent scheduling delays, as CL jobs must wait for enough devices to become available before training can begin.
Device heterogeneity. Edge devices in CL vary widely in hardware capabilities, including memory and CPU capacity, as well as in data availability and software version, leading to inconsistent response times. Figure 2(b) showcases this heterogeneity, focusing on variations in memory and CPU capacities among edge devices, based on data from AI Benchmark Ignatov et al. (2019). It also annotates the minimum hardware requirements needed to execute three popular on-device ML models within a reasonable time. Additionally, CL jobs often target specific device subsets based on factors like data availability, software version, or hardware capacity. These subsets can overlap, nest, or be mutually exclusive, leading to complex resource contention patterns where jobs compete for the same devices.
Together, dynamic availability and device heterogeneity slow down CL training, prompting solutions like client selection Lai et al. (2021b); Abdelmoniem et al. (2023); Balakrishnan et al. (2022), quantization Reisizadeh et al. (2020), and model aggregation He et al. (2023) to optimize single-job performance.
@@
(b) Random Matching ($\bar{JCT}$=12).
(c) Shortest remaining service first (SRSF) ($\bar{JCT}$=11).
(d) Optimal ($\bar{JCT}$=9.3).
Figure 3: Toy example showing three resource schedules across multiple CL jobs. Job demands and resource eligibility are shown in the top row. Devices check in at a constant rate. Eligible devices only for Emoji jobs are marked with blue; all devices are eligible for the Keyboard job. The label of each client indicates its job assignment. Random Matching and SRSF inefficiently allocate scarce Emoji-eligible devices to Job 1, which already has sufficient Keyboard-eligible resources. In contrast, the optimal schedule allocates these scarce resources to Job 2 followed by Job 3, minimizing the average JCT.
2.2 Multiple CL Jobs’ Resource Management
To orchestrate multiple CL jobs, several CL resource managers have been proposed with three primary designs.
@@ 2.2 Multiple CL Jobs’ Resource Management
    Apple’s CL resource management Paulik et al. (2021) is driven by clients, where each client independently samples from a list of CL jobs they are able to execute.
@@
    Meta’s CL resource manager Huba et al. (2022) is centralized, where it randomly matches each client with one eligible CL job.
@@
    Google’s CL resource manager Bonawitz et al. (2019) is driven by jobs, where each job independently samples from available clients.
Despite the seeming variety in their designs, existing resource managers all rely on random device-to-job matching. This approach works when devices are abundant but fails to address resource contention when demand exceeds supply, resulting in longer scheduling delays and inefficient resource use.
2.3 Limitation and Opportunities
@@
To highlight the shortcomings of existing methods, consider the toy example in Figure 3, which compares three scheduling strategies for three CL jobs—“Emoji Prediction Job 1,” “Emoji Job 2,” and a “Keyboard Prediction Job”—competing for devices with varying eligibility. We assume all jobs arrive simultaneously and clients with different eligibilities become online constantly over time. Both random matching and SRSF waste scarce resources (i.e., Emoji clients) on the Keyboard job, which has ample device options, while the optimal schedule prioritizes efficiency, reducing JCT to 9.3 time units. Note that the contention patterns in real-world scenarios are often more complex and larger in scale than the one presented in this example.
@@
Impact of resource contention.
Single-job FL optimizations, such as client selection Lai et al. (2021b); Abdelmoniem et al. (2023); Balakrishnan et al. (2022), often assume that CL jobs have access to sufficient number of online devices at any time. However, in practice, there could be multiple CL jobs running at the same time and competing for the same set of devices, leading to resource contention Bonawitz et al. (2019). This can significantly impact the performance of CL jobs, such as accuracy and end-to-end training time.
We analyze the impact of resource contention on CL jobs’ performance in Figure 5. In this experiment, the resource pool is evenly partitioned and managed by each job, who aims to train a ResNet-18 model with 100 clients per round on the FEMNIST dataset Cohen et al. (2017). We vary the number of concurrently running jobs to observe its impact. As more jobs share the same device pool, the available device choices for each job become increasingly constrained, leading to a noticeable degradation in the round-to-accuracy performance. Hence, evenly partitioning resource pool for each CL job is no better than a shared device pool in terms of participant diversity.
Breakdown of request completion time.
While most existing CL efforts, such as quantization Reisizadeh et al. (2020) and client selection He et al. (2023), predominantly focus on model convergence rate or response collection time—–i.e., the time needed to collect a sufficient number of responses—–they often overlook a critical component: scheduling delay, as depicted in Figure 1.
Figure 5 breaks down JCT for a single training round under varying contention levels, using the same setup as Figure 5. Utilizing the same experimental setup as in Figure 5, we quantify both the average scheduling delay and response collection time with random device-to-job matching during one round of training and testing. The shaded regions cover the duration for each individual job. As our results in Figure 5 indicate, scheduling delay can significantly impact overall JCT, especially when resource supply falls short of demand.
@@ 3 Venn Overview
Venn serves as a standalone CL resource manager that operates at a layer above all CL jobs, and it is responsible for allocating each checked-in resource to individual jobs. Figure 6 illustrates Venn’s workflow and its role within the lifecycle of an CL job.
@@
In each execution round, a job submits resource requests to Venn, specifying its device requirements (e.g., minimum CPU capacity) and resource demands (the number of devices needed per round) (⓪). Meanwhile, edge devices continuously check in with Venn as they become available over time (①). Based on the real-time resource demand and supply, Venn generates a resource allocation plan to assign one CL job to each checked-in device until the job’s needs are met (②). Upon receiving the task assignment from Venn, each device adheres to the allocation plan and participates in the corresponding job (③). The device then retrieves the computation plan from the job and performs on-device training or inference (Lv et al., 2022; CoreML, ) (③ and ④). Finally, devices either submit their training results to the job or disconnect if their availability changes (e.g., battery runs low) (⑤).
Steps ③ to ⑤ follow standard CL protocols between jobs and devices, while Venn focuses on optimizing the job-to-device assignment in Step ②. This allocation step is critical for managing resource contention and reducing scheduling delays. A detailed breakdown of responsibilities is provided in Appendix A.